# On the Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ over Stable Semirings

### v1

@@ 7 Related Work
Computing the least fixpoint solution to a recursive $\text{\sf Datalog}^{\circ}$ program boils down to solving fixpoint equations over semirings. In particular, we are given a multi-valued polynomial function $f:S^{n}\to S^{n}$ over a commutative semiring, and the problem is to compute a (pre-) fixpoint of $f$, i.e. a point $x\in S^{n}$ where $x=f(x)$. As surveyed in much details in [13], this problem was studied in a very wide range of communities, such as in automata theory [16], program analysis [5, 20], graph algorithms [4, 19, 18] communities since the 1970s. (See [23, 11, 17, 10, 28] and references thereof).
@@
When $f=Ax+b$ is linear, as shown in the paper $f^{(k)}(x)=A^{(k-1)}b$ and thus at fixpoint the solution is $A^{(\omega)}b=\lim_{k\to\infty}A^{(k-1})b$, interpreted as a formal power series over the semiring. If there is a finite $k$ for which $A^{(k)}=A^{(k+1)}$, then it is easy to see that $A^{(\omega)}=A^{(k)}$. The problem of computing $A^{(\omega)}$ is called the algebraic path problem [23], which unifies many problems such as transitive closure [26], shortest paths [8], Kleene’s theorem on finite automata and regular languages [14], continuous dataflow [5, 12], etc. If $A$ is a real matrix, then $A^{(\omega)}=I+A+A^{2}+\cdots$ is exactly $(I-A)^{-1}$, if it exists [3, 24, 9].
@@
There are several classes of solutions to the algebraic path problem, which have pros and cons depending on what we can assume about the underlying semiring (whether or not there is a closure operator, idempotency, natural orderability, etc.). We refer the reader to [10, 23] for more detailed discussions. Here, we briefly summarize the main approaches.
The first approach is to keep iterating until a fixpoint is reached; in different contexts, this has different names: the naive algorithm, Jacobi iteration, Gauss-Seidel iteration, or Kleene iteration. The main advantage of this approach, which is studied in this paper, is that it assumes less about the underlying algebraic structure: we do not need both left and right distributive law, and do not need to assume a closure operator. The second approach is based on Gaussian elimination (Gauss-Jordan elimination), which, assuming oracle access to the solution $x^{*}$ of the 1D problem $x=1+ax$, can solve the algebraic path problem in $O(n^{3})$-time [23, 17]. The third approach is based on specifying the solutions based on the free semiring generated when viewing $\bm{A}$ as the adjacency matrix of a graph [25]. The underlying graph structure (such as planarity) may be exploited for very efficient algorithms [19, 18].
We also discuss the previous work on the fixed point solution to $x=f(x)$ in the non-linear case. See [22, 21, 15] for the study when $f$’s component functions are multivariate polynomials over the Kleene algebra. In the non-linear case, Kleene iteration may not converge in a finite number of steps. Thus, methods based on Galois connection or on widening/narrowing approaches [6] were proposed. These approaches are (discrete) lattice-theoretic. More recently, a completely different approach drawing inspiration from Newton’s method for solving a system of (real) equations was proposed [11, 7]. Newton’s method converges at least as fast as Kleene iteration, but each iteration could be more costly it need to compute a matrix inverse [7]. Recently, semi-naïve evaluation has been extended for a higher-order functional language called Datafun [2]. The book [10] contains many fundamental results on algebraic structures related to semirings and computational problems on such structures.
@@ Appendix A Convergence of Naturally Ordered Semirings
Proof.
Consider $A^{(k)}x$ as $k$ increases for any fixed $x$. If there is a $k\leq nL$ such that $A^{(k)}x=A^{(k+1)}x$ then convergence has been reached within the desired number of steps. Otherwise when $A^{(k)}x\neq A^{(k+1)}x$ there exists an $i$ such that dimension $i$ in $A^{(k+1)}x$ is strictly greater than dimension $i$ in $A^{(k)}x$. This can only occur $L$ times for each $i$ by definition of the partial order. Knowing that there are at most $n$ dimensions in $A^{(k)}x$, the lemma follows. ∎
@@
Consider the following graph corresponding to a $n\times n$ matrix $A$. There is a cycle on $n$ nodes. Order the edges from $1$ to $n$. Each edge is labeled 00 except the edge from $1$ to $2$, which is labeled as $1$. Traversing the cycle $k$ times and multiplying the labels of the edges returns $\min\{k,L\}$. It takes a walk of length $\Omega(nL)$ to reach the element $L$. ∎
Appendix B Missing Proof from Section 4.1
We show that the semiring we defined in Section 4.1 is indeed a (commutative) semiring.
Monoid $(S,+)$ with identity $\mathcal{O}$:
  • $A+\mathcal{O}=\mathcal{O}+A=A$. This follows from the definition.
  • $(A+B)+C=A+(B+C)$. When $A,B,C\neq\mathcal{O}$, this is equivalent to showing $\max\{\max\{A_{i},B_{i}\},C_{i}\}=\max\{A_{i},\max\{B_{i},C_{i}\}\}$ for all $i\in[m]$, which follows from $\max$ being commutative. If $A,B$, or $C$ is $\mathcal{O}$, then it is easy to check that it holds true.
Monoid $(S,\cdot)$ with identity $\emptyset$:
  • $A\cdot\emptyset=\emptyset\cdot A=A$. If $A\neq\mathcal{O}$, this is immediate from the definition. If $A=\mathcal{O}$, again by definition, $\mathcal{O}\cdot\emptyset=\mathcal{O}\cdot\emptyset=\mathcal{O}$.
  • $(A\cdot B)\cdot C=A\cdot(B\cdot C)$. When $A,B,C\neq\mathcal{O}$, it is an easy exercise to see $((A\cdot B)\cdot C)_{i}=(A\cdot(B\cdot C))_{i}=\min\{A_{i}+B_{i}+C_{i},p\}$. If $A,B$ or $C$ is $\mathcal{O}$, both sides become $\mathcal{O}$.
Commutative:
  • $A+B=B+A$. If $A,B\neq\mathcal{O}$, we have $(A+B)_{i}=\max\{A_{i},B_{i}\}=(B+A)_{i}$. Otherwise, it is immediate from the definition of $\mathcal{O}$.
  • $A\cdot B=B\cdot A$. We also show that the multiplication is also commutative. If $A,B\neq\mathcal{O}$, we have $(A\cdot B)_{i}=\min\{A_{i}+B_{i},p\}=\min\{B_{i}+A_{i},p\}=(B\cdot A)_{i}$. Otherwise $A\cdot B=B\cdot A=\mathcal{O}$ from the definition.
$\mathcal{O}$ is an multiplicative annihilator: We have $\mathcal{O}\cdot A=A\cdot\mathcal{O}=\mathcal{O}$ for all $A\in S$ from the definition.
Distributive:
  • $A\cdot(B+C)=A\cdot B+A\cdot C$. Assume that $A,B,C\neq\mathcal{O}$ since otherwise it is straightforward to see that it holds true. We then have,
    $(A\cdot(B+C))_{i}$ $=\min\{A_{i}+\max\{B_{i},C_{i}\},p\}=\min\{\max\{A_{i}+B_{i},A_{i}+C_{i}\},p\}$
    $=\max\{\min\{A_{i}+B_{i},p\},\min\{A_{i}+C_{i},p\}\}=\max\{(A\cdot B)_{i},(A\cdot C)_{i}\}$
    $=(A\cdot B+A\cdot C)_{i}$
  • $(B+C)\cdot A=B\cdot A+C\cdot A$. The proof is symmetric.
Appendix C Warm-up for Proof of Theorem 1.8
@@
For some $T$ (possibly empty), we have $W=X_{1}X_{2}X_{3}\ldots X_{p+1}T$. By definition $\Phi(X_{1})=\Phi(X_{1}X_{2})=\ldots=\Phi(X_{1}X_{2}\ldots X_{p+1})$, we have
$\Phi(W)$ $=\Phi(X_{1}X_{2}X_{3}\ldots X_{p+1}T)$
$=\Phi(X_{1}X_{2}X_{3}\ldots X_{p}T)$
[…]
$=\Phi(X_{1}T),$
Here, we used the fact that $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ all are walks from $i$ to $j$ since $X_{1}$, $X_{1}X_{2}$, …, $X_{1}X_{2}\ldots X_{p+1}$ all end at the point where $T$ starts. Further, all walks $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ are strictly shorter than $W$. This implies that element $\Phi(W)$ appears at least $p$ times in $A^{(k)}_{i,j}$. Thanks to Proposition 2.2, we conclude $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired. ∎
@@
Consider a walk $W^{\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup(B\setminus A)\cup([H]\setminus(A\cup B))=[H]\setminus(A\setminus B)$ and ends with $T^{\prime}$. Similarly, consider a walk $W^{\prime\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup([H]\setminus(A\cup B))$ and ends with $T^{\prime}$. Note that $W$ and $W^{\prime}$ are different since $B\setminus A\neq\emptyset$. Further they are walks from $i$ to $j$ since every cycle $C_{h}$, $h\in[H]$ starts from and ends at the same vertex $v$. Further, both walks are shorter than $W$, and therefore are in $\mathcal{W}^{(k)}_{i,j}$. Since we have
$\Phi(W^{\prime})$ $=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}(B\setminus A)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$
$\Phi(W^{\prime\prime})$ $=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$
We will show using 1-stability of the semiring that $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$, implying $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Thus, it suffices to show $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$. To see this note the following:

### v2

@@ 6 Related Work
If the semiring is naturally ordered222 $\bm{S}$ is naturally ordered if the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$ is a partial order. , then the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program is the least fixed point of $f$ under the same partial order extended to $S^{n}$ componentwise. This is the least fixpoint semantics of a $\text{\sf Datalog}^{\circ}$ program. The naïve evaluation algorithm for evaluating Datalog programs extends naturally to evaluating $\text{\sf Datalog}^{\circ}$ programs: starting from $\bm{x}=0^{n}$, we repeatedly apply $f$ to $\bm{x}$ until a fixpoint is reached $\bm{x}=f(\bm{x})$. The core semiring of a POPS is naturally ordered. Thus, we can find the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program by applying the naïve evaluation algorithm [12].
@@
Computing the least fixpoint solution to a recursive $\text{\sf Datalog}^{\circ}$ program boils down to solving fixpoint equations over semirings. In particular, we are given a multi-valued polynomial function $f:S^{n}\to S^{n}$ over a commutative semiring, and the problem is to compute a (pre-) fixpoint of $f$, i.e. a point $x\in S^{n}$ where $x=f(x)$. As surveyed in in [12], this problem was studied in a very wide range of communities, such as in automata theory [14], program analysis [4, 18], and graph algorithms [3, 17, 16] since the 1970s. (See [19, 9, 15, 8, 23] and references thereof).
@@
When $f=Ax+b$ is linear, as shown in the paper $f^{(k)}(x)=A^{(k-1)}b$ and thus at fixpoint the solution is $A^{(\omega)}b=\lim_{k\to\infty}A^{(k-1)}b$, interpreted as a formal power series over the semiring. If there is a finite $k$ for which $A^{(k)}=A^{(k+1)}$, then it is easy to see that $A^{(\omega)}=A^{(k)}$. The problem of computing $A^{(\omega)}$ is called the algebraic path problem [19], which unifies many problems such as transitive closure [21], shortest paths [5], Kleene’s theorem on finite automata and regular languages [13], and continuous dataflow [4, 11]. If $A$ is a real matrix, then $A^{(\omega)}=I+A+A^{2}+\cdots$ is exactly $(I-A)^{-1}$, if it exists [2, 20, 7].
@@
There are several classes of solutions to the algebraic path problem, which have pros and cons depending on what we can assume about the underlying semiring (whether or not there is a closure operator, idempotency, natural orderability, etc.). We refer the reader to [8, 19] for more detailed discussions.
@@ Appendix A Convergence of Naturally Ordered Semirings
Proof A.5.
Consider $A^{(k)}x$ as $k$ increases for any fixed $x$. If there is a $k\leq nL$ such that $A^{(k)}x=A^{(k+1)}x$ then convergence has been reached within the desired number of steps. Otherwise when $A^{(k)}x\neq A^{(k+1)}x$ there exists an $i$ such that dimension $i$ in $A^{(k+1)}x$ is strictly greater than dimension $i$ in $A^{(k)}x$. This can only occur $L$ times for each $i$ by definition of the partial order. Knowing that there are at most $n$ dimensions in $A^{(k)}x$, the lemma follows.
@@
Consider the following graph corresponding to a $n\times n$ matrix $A$. There is a cycle on $n$ nodes. Order the edges from $1$ to $n$. Each edge is labeled 00 except the edge from $1$ to $2$, which is labeled as $1$. Traversing the cycle $k$ times and multiplying the labels of the edges returns $\min\{k,L\}$. It takes a walk of length $\Omega(nL)$ to reach the element $L$.
Appendix B Warm-up for Proof of Theorem 1.5
@@
For some $T$ (possibly empty), we have $W=X_{1}X_{2}X_{3}\ldots X_{p+1}T$. By definition $\Phi(X_{1})=\Phi(X_{1}X_{2})=\ldots=\Phi(X_{1}X_{2}\ldots X_{p+1})$.
Thus, $\Phi(W)=\Phi(X_{1}X_{2}X_{3}\ldots X_{p+1}T)$ $=\Phi(X_{1}X_{2}X_{3}\ldots X_{p}T)$ $\ldots=\Phi(X_{1}T)$, This uses the fact that $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ all are walks from $i$ to $j$ since $X_{1}$, $X_{1}X_{2}$, …, $X_{1}X_{2}\ldots X_{p+1}$ all end where $T$ starts. Further, all walks $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ are strictly shorter than $W$. This implies that $\Phi(W)$ appears at least $p$ times in $A^{(k)}_{i,j}$. Using Proposition 2.2, we conclude $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
@@
Consider a walk $W^{\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup(B\setminus A)\cup([H]\setminus(A\cup B))=[H]\setminus(A\setminus B)$ and ends with $T^{\prime}$. Similarly, consider a walk $W^{\prime\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup([H]\setminus(A\cup B))$ and ends with $T^{\prime}$. Note that $W$ and $W^{\prime}$ are different since $B\setminus A\neq\emptyset$. Further they are walks from $i$ to $j$ since every cycle $C_{h}$, $h\in[H]$ starts from and ends at the same vertex $v$. Further, both walks are shorter than $W$, and therefore are in $\mathcal{W}^{(k)}_{i,j}$. Since we have $\Phi(W^{\prime})=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}(B\setminus A)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$ and $\Phi(W^{\prime\prime})=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$. We will show using 1-stability of the semiring that $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$, implying $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Thus, it suffices to show $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$. To see this: